### v1

# Semantic Component Analysis: Discovering Patterns in Short Texts Beyond Topics

Abstract
Topic modeling is a key method in text analysis, but existing approaches are limited by assuming one topic per document or fail to scale efficiently for large, noisy datasets of short texts. We introduce Semantic Component Analysis (SCA), a novel topic modeling technique that overcomes these limitations by discovering multiple, nuanced semantic components beyond a single topic in short texts which we accomplish by introducing a decomposition step to the clustering-based topic modeling framework. Evaluated on multiple Twitter datasets, SCA matches the state-of-the-art method BERTopic in coherence and diversity, while uncovering at least double the semantic components and maintaining a noise rate close to zero while staying scalable and effective across languages, including an underrepresented one.
1 Introduction
Topic modeling is an essential tool for automated text analysis in digital humanities, computational social science, and related disciplines Grimmer et al. (2022); Rauchfleisch et al. (2023); Blei (2012); Maurer and Nuernbergk (2024) where it aims to find high-level semantic patterns, the topics, in a corpus and then assigning them to documents. The state-of-the-art for large-scale datasets, BERTopic Grootendorst (2022), assumes a single topic per document, which often does not hold even for short texts Paul and Girju (2010), as can be seen in Figure 1. BERTopic and related approaches do thus not model a large portion of the information in the documents. Besides, while they are able to achieve high topic coherence and diversity, a high rate of samples is usually assigned to no topic, i.e. they have a high noise rate Egger and Yu (2022).
Figure 1: A Tweet from the Trump dataset decomposed into its semantic components as defined by the Top-3 component activation scores of iteration 1 (components 6, 24, and 0) and subsequent iterations (components 56, 118, 70). Component 0 refers to the most dominant topic in the Trump dataset, of which a majority of Tweets is about Donald Trump himself.
To overcome these limitations, we propose Semantic Component Analysis (SCA) as a novel method to discover multiple patterns per document, which we coin “semantic components". SCA is a topic modeling technique, that - in addition to topics - formalizes the discovery of nuanced semantic components and scales to large datasets while maintaining a noise rate close to zero and matching BERTopic in topic coherence and diversity. Based on a sequence embeddings-based clustering pipeline as proposed by Angelov (2020), for each cluster, we introduce a decomposition step and iterate over the residual embeddings. This decomposition enables us to represent a sample’s content independently of the semantic components found in previous iterations and to find additional components.
Figure 2: Illustration of the basic clustering algorithm for linear feature discovery. Top-Right: Top-4 components (by cluster size) of the first iteration, that is equivalent to a BERTopic run. Bottom-right: components uncovered by SCA in subsequent iterations (see Table 7 in the appendix for more examples).
We evaluate our method on three Twitter datasets, one of all the Tweets of Donald Trump (Brown, 2021), another one compiled of Chinese-language news outlets that we scraped ourselves and publish alongside this paper, and finally a corpus of Tweets in Hausa (Muhammad et al., 2022). We show that our method is able to match the performance of BERTopic (Grootendorst, 2022) in terms of topic coherence and diversity, while additionally providing a large number of semantic components missed by the state-of-the-art approach. We also show that our method can provide an interpretable transformation that matches the performance of the base embedding in a variety of downstream tasks.
Our main contributions are as follows:
  • We propose a novel topic modeling method, Semantic Component Analysis (SCA), that can discover multiple semantic components per document.
  • We show that our method is able to match state-of-the-art topic modeling, while providing additional, more nuanced insights into the semantic patterns in the data and maintaining a noise rate close to zero.
  • We publish a dataset of Tweets from Chinese news outlets and use a dataset of Tweets in Hausa, a low-resource language, to show the generalizability of our method beyond English.
2 Related Work
Topic modeling’s goal is to find semantic patterns in a corpus and to assign them to the documents. Latent Dirichlet Allocation (LDA) (Blei et al., 2003; Pritchard et al., 2000) works well on small datasets of long texts and has been generalized to larger datasets Hoffman et al. (2012). Modeling a distribution over topics for each document, it can account for multi-topic documents. However, it performs poorly on short texts Ushio et al. (2021); de Groot et al. (2022), is generally vulnerable to noisy data, and produces topics that lack domain adaption Egger and Yu (2022). Recent advancements have integrated neural networks (Srivastava and Sutton, 2017; Liu et al., 2015; Dieng et al., 2019; Qiang et al., 2016; Zhao et al., 2021). Topic modeling based on LLMs Pham et al. (2024); Doi et al. (2024) has also gained attention, though their cost is a barrier for most large-scale applications. Methods based on clustering embeddings stand out because of their efficiency, domain-adaption, and topic coherence on large corpora of short texts. Top2Vec (Angelov, 2020) clusters document embeddings generated by Doc2Vec (Le and Mikolov, 2014). There, UMAP (McInnes et al., 2020) reduces embedding dimensionality mitigating the “curse of dimensionality" (Reimers and Gurevych, 2021), and clustering is performed using HDBSCAN (Campello et al., 2015). BERTopic (Grootendorst, 2022) improves the Top2Vec framework by using SBERT embeddings (Reimers and Gurevych, 2019) and c-TF-IDF weighting for representing the topics. It is considered state-of-the-art for large datasets of short texts Egger and Yu (2022). However, BERTopic does not provide a way to generalize to multi-topic documents. The cluster-hierarchy of HDBSCAN could be used to assign multiple topics per sample, but this approach is insufficient for finding patterns that are not subtopics of each other. Besides, BERTopic’s results usually suffer from a high noise rate Egger and Yu (2022), which can range at up to 74% de Groot et al. (2022). In summary, the topic modeling field does not provide a solution for scenarios with large, noisy datasets of short texts containing multiple patterns of interest per sample, that matches the state-of-the-art topic quality.
3 Methodology
SCA is a clustering pipeline that introduces a novel decomposition step to find semantic components in the embedding space, which are represented by unit vectors. For clustering, we employ the cosine distance in UMAP to map the embeddings to a 5-dimensional, Euclidean space on which HDBSCAN operates, mirroring the setup in BERTopic. Visualized in Figure 2, it works as follows:
  • 0.
    Embed input sequences
  • 1.
    Reduce and Cluster the embeddings $\mathbf{x}_{j}$
  • 2.
    Represent component defined by cluster $C_{i}$ through normalized centroid $\mathbf{v}_{i}\in\mathbb{R}^{D}$
    $\mathbf{v}_{i}=\frac{\mathbf{v}^{\prime}_{i}}{||\mathbf{v}^{\prime}_{i}||},\ \ \mathbf{v}^{\prime}_{i}=\frac{1}{|C_{i}|}\sum_{\mathbf{x}_{j}\in C_{i}}\mathbf{x}_{j}$ (1)
  • 3.
    Decompose the $\mathbf{x}_{j}$ into their linear components along each $\mathbf{v}_{i}$ and the residual $\mathbf{x}_{i}^{\prime}$:
    $\mathbf{x}_{i}^{\prime}=\mathbf{x}_{i}-\mu\mathbb{1}_{\alpha_{i,j}>\alpha}\langle\mathbf{x}_{j},\mathbf{v}_{i}\rangle\mathbf{v}_{i}$ (2)
    where $\alpha_{i,j}=\frac{\langle\mathbf{x}_{j},\mathbf{v}_{i}\rangle}{||\mathbf{x}_{j}||}$
  • 4.
    Repeat steps 1-3 on residual embeddings
  • 5.
    Represent and Merge components by c-TF-IDF weighting and token overlap
After each decomposition step in 3, we set $\mathbf{x}_{j}=\mathbf{x}_{j}^{\prime}$ for all $j$ and repeat it for the next cluster. The algorithm converges when the residuals are zero, but we introduce other practical stopping criteria listed in Appendix 8.3. Note, that we introduce two hyperparameters: $\mu\in[0,1]$ controlling the amount of the decomposition and a threshold $\alpha$, which is compared against $\alpha_{i,j}$, the cosine similarity between $\mathbf{x}_{j}$ and $\mathbf{v}_{i}$. See Appendix 8.1 for a detailed discussion of $\alpha$ and $\mu$. We assume that each cluster found represents a semantic component. To obtain an interpretable representation, we follow BERTopic and use class-based (c-) TF-IDF to rank tokens according to their importance for a cluster, from which we obtain the TOP-10 set of tokens for each component.
We introduce a merging step that computes the token overlap between two components as the fraction of tokens their representations share. Here, we introduce a threshold $\theta\in[0,1]$ and merge two components into a multi-dimensional component if the token overlap is larger than $\theta$ and keep the representation of the component with the lower ID (i.e. the component that appeared earlier in the procedure). That way, we avoid duplicates arising in larger clusters spread across a non-linear or multi-dimensional part of the embedding space.
Finally, we include the Medoid of each cluster, which is the sample closest to the centroid in terms of cosine similarity. We note, that SCA employs the same modular approach as BERTopic, which allows the embedding, dimensionality reduction, clustering, and representation method to be replaced by similar methods.111see https://maartengr.github.io/BERTopic/ Furthermore, our specific choices for these parts of the model, namely UMAP and HDBSCAN, make the first iteration of SCA equivalent to a BERTopic run, which allows for a direct comparison of the methods.
3.1 Evaluation
We evaluate our method on three datasets of different sizes and languages. [^2] https://github.com/mainlp/semantic_components For an English-language use-case, we include all Tweets by Donald Trump until December 2024 (60K samples, Brown (2021)). To evaluate on a high-resource non-English language, we collected and publish 1.5M Chinese-language Tweets from a list of news outlets (details in Appendix 9.1). To cover a low-resource language, we include a set of 512K Tweets in Hausa collected by Muhammad et al. (2022). We evaluate our method in terms of number of components, noise rate, topic coherence, and diversity (see Appendix 9.2 for definitions) and compare it to the highly popular BERTopic. Furthermore, we perform a grid run over the hyperparameters to assess their influence. To check if the components found are indeed beyond what is found by BERTopic, we compute the token overlap between the components found by SCA and the topics found by BERTopic, as well as the sample overlap with the cluster hierarchy in HDBSCAN of BERTopic and the components found by SCA (see Appendix 9.2 for details). Finally, we evaluate the transformation of the embeddings into the component space in terms of its performance in downstream tasks in the MTEB benchmark (Muennighoff et al., 2023).
4 Results
Statistic Trump Hausa Ch. News
BERTopic SCA BERTopic SCA BERTopic SCA
Number of Components $\uparrow$ 55 182 102 331 212 594
Noise Rate $\downarrow$ 0.275 0.000 0.574 0.001 0.429 0.000
NPMI Coherence $\uparrow$ -0.135 -0.168 -0.028 -0.061 0.186 0.120
CV Coherence $\uparrow$ 0.370 0.360 0.431 0.400 0.683 0.595
Topic Div. $\uparrow$ 0.947 0.703 0.909 0.804 0.929 0.803
Avg. Max. Sample Overlap $\in[0,1]$downarrow 0.079 0.093 0.077
Avg. Max. Token Overlap $\in[0,10]$downarrow 1.387 1.852 1.318
[…]
We provide a table of the Top-10 components of the first three iterations of runs on each dataset in Appendix 9. A subset can be found in Figure 2. The metrics table can be found in Figure 3 while that of the grid runs can be found in Appendix 9.2. For all datasets, the method finds a large number of components, with the Trump dataset yielding 182 components in total after merging, the Chinese News dataset 594 and the Hausa dataset 331, surpassing the 55, 212, and 102 respective topics found by BERTopic. These components relate to patterns distinct from the topics found by BERTopic, which is evident from the low average maximum token overlap scores of 1.85, 1.39 and 1.32 on the three datasets respectively. On the Trump dataset, only 8 components with a token overlap of higher than 0.5 are merged, on the Hausa dataset it is 9, and for the Chinese News dataset 19. The average maximum sample overlaps with any of BERTopic’s hierarchical subtopics are small and between 0.077 and 0.093 on all datasets. This indicates that the additional components are not sub- or supertopics of the topics found by BERTopic. The noise rate is close to zero for all datasets, i.e. almost all samples are assigned at least one component. The topic coherence scores are high for all datasets, with the Chinese dataset showing the highest NPMI coherence of 0.120 and the Trump dataset showing the lowest of -0.168, all matching the baseline. The topic diversity is high for all datasets, between 0.7 and 0.8, but about 10-15% lower than in the baseline. In sum, this shows that the quality of the components found matches that of BERTopic while providing a more nuanced view of the data through the large number of extra components found and the low noise rate. Comparing the results over the hyperparameters on the Trump dataset, we find that for $\alpha>0.7$ and $\mu<0.4$, the method reliably finds a large number of semantic components at least double the number of topics found by BERTopic, usually ranging from three to four times.
The additional components found relate to patterns that intuitively match the expected distribution of the datasets. For example, in the Trump dataset, beyond the topics about Trump himself, entrepreneurship, and election campaigning found by BERTopic, we uncover components about “fake news", inspiration and role models, paying condolences, and the “witch hunt" narrative among others (see Appendix 9.3) underlining SCA’s ability to find important patterns not captured by the baseline. This view is reinforced by the component distributions of single samples, as shown in Figure 1. Similarly, in the Chinese dataset, the additional components relate to important motives in Chinese news reporting, such as protests, family, the trade war with the US, and currency (see Appendix 9.4). In the Hausa dataset, the additional insights include components about poverty, courts and law, and work (see Appendix 9.5). However, we note, that in this low-resource scenario, some components relate to semantically more narrow concepts, which is possibly due to a lack of expressive power in the base embeddings.
Finally, the results of the topic distribution provided by SCA can match the performance of PCA and the base embedding model in the MTEB benchmark (see Appendix 9.6), indicating the method’s faithfulness in representing the information of the samples in the component space.
5 Conclusion
We introduce Semantic Component Analysis as a novel topic modeling method that is able to discover multiple semantic components per document. We show, that the components found by SCA indeed provide new, more nuanced insights into the data while maintaining a noise rate close to zero and matching the state-of-the-art in terms of topic quality. Furthermore, we show that the method is able to generalize to multiple languages, including a low-resource African language, and provide a faithful transformation of the embeddings into the component space. We believe that SCA provides a significant improvement in large-scale text analysis and opens up new possibilities for the interpretation of semantic patterns in short texts.
6 Limitations
While SCA provides a novel perspective on topic modeling by providing an efficient way to discover multiple, nuanced semantic components per sample and thus a more detailed explanation of the pattern distribution in a dataset, it has its limitations. The method is not able to capture non-linear, multi-dimensional components in the data, as it is based on a linear decomposition of the embedding space. In other words, SCA is based in the linear representation hypothesis Park et al. (2024). To some extent, the iterative procedure of SCA is able to mitigate this problem: Because each cluster is represented by the centroid, the decomposition might retain a part of the cluster-information in the residual embeddings in cases, when one linear unit vector is not able to capture the full complexity in terms of dimension or linearity. In such cases, similar components might be found in subsequent iterations, which can be merged by the overlap threshold.
Furthermore, while we show SCA’s applicability to a range of datasets and languages, we restrict our investigation to short texts, more precisely Tweets. We argue that problems with longer documents can partially be mitigated by splitting them into smaller parts, such as sentence level, and apply SCA to the resulting dataset. Since SCA scales to large numbers of samples even on small systems, this should be feasible for most applications.
As is common with topic modeling, the method’s token-based representation of topics has its limitations, too, as it might not able to capture the full complexity of the semantic components found. While additional methods like LLM summarizations and Medoid representations can provide additional insights, a core problem with more expressive approaches are uncertain faithfulness, a lack of good evaluation metrics as well as increased computational cost.
Acknowledgments
We’re grateful to Kristin Shi-Kupfer and David Adelani for consulting on the Chinese and Hausa datasets respectively. Furthermore, we would like to thank Xinpeng Wang, Ryan Soh-Eun Shim, Beiduo Chen, and Yang Janet Liu for their valuable feedback on earlier drafts of this paper.
Appendix
7 Risks and Ethical Considerations
There are several risks and ethical considerations associated with the use of SCA.
Firstly, there is a risk of misinterpretation or misrepresentation of the discovered semantic components. The interpretation of these components relies on the assumption that they represent meaningful and coherent topics or themes in the data. However, there is always a possibility of misinterpreting or misrepresenting the true meaning of these components, leading to incorrect conclusions or biased interpretations. Even earlier in the pipeline, there is a risk of algorithmic bias in the clustering and representation process. Especially the base embedding model can introduce biases based on the training data and the model architecture that SCA can inherit. It is important to carefully select and validate thesethe base model to ensure as fair and unbiased representations as possible.
Secondly, there is a risk of privacy infringement when working with sensitive or personal data. SCA relies on analyzing and clustering text data, which may contain sensitive information about individuals or groups. It is important to handle and protect this data in accordance with privacy regulations and ethical guidelines and to ensure that the semantic components extracted from the data do not reveal sensitive or personally identifiable information to unauthorized parties. The usage of the data in our evaluation does not expose any individual to such risks, as the data was either publicly available to begin with (the Hausa Dataset and Trump Archive) or does not contain any data produced by private Twitter accounts, as the Chinese News Data we scrape only consists of Tweets by the official accounts of the news outlets we include.
Thirdly, the nature of online social media data, such as Tweets, can introduce additional risks and challenges. Social media data can be subject to manipulation or misinformation, and contain offensive or harmful content. SCA may inadvertently amplify or propagate such content if not properly filtered or evaluated. It is important to consider the potential impact of the data and the extracted semantic components on individuals, communities, and society as a whole. Furthermore, the data we collect and
Lastly, there may be ethical considerations regarding the use of SCA in certain domains or applications. For example, if SCA is used in the context of social media analysis or opinion mining, there is a risk of amplifying or reinforcing existing biases or stereotypes present in the data. It is important to critically evaluate and address these ethical considerations when applying SCA in such contexts.
Overall, it is crucial to approach the use of SCA with caution, considering the potential risks and ethical implications associated with its application.
8 Methodology Appendix
8.1 Hyperparameter Discussion
Our method involves two hyperparameters that influence the decomposition step, $\alpha$ and $\mu$. Intuitively, both can be used to influence the level of superposition allowed in the decomposition, i.e. the number of semantic components that can occupy linearly dependent parts of the embedding space and thus the the level of detail and the number of discovered semantic components. This can be necessary in cases of large data where the number of semantic components assumed is higher than the dimension of the embedding space.
In the simplest case, when no superposition is assumed, we set $\mu=1$ and $\alpha=0$. We illustrate this version of the algorithm in Figure 2. Setting $\mu<1$ is the first strategy to allow for superposition. This way, each decomposition step will only remove a fraction of the feature from the activation, allowing subsequent clustering passes to discover more, linearly dependent features. The second strategy is to set $\alpha>0$, which introduces conditional decomposition. Here, a feature is only decomposed from an activation if the cosine similarity between the activation and the feature is above the threshold $\alpha$.
Furthermore, we introduce the merging threshold $\theta$ to merge components that are similar in terms of their token representation. This is necessary to avoid duplicate components that arise for large clusters spread across a non-linear or multi-dimensional part of the embedding space. We note, that the choice of $\theta$ is somewhat arbitrary and should be chosen based on the specific use case. In our evaluation, we set $\theta=0.6$ for the Trump dataset, which resulted in 11 components being merged. For the Chinese and Hausa datasets, we did not employ component merging.
8.2 SCA as a Transformation
The decomposition step in SCA is a linear transformation of the embeddings into a space where each dimension corresponds to the score along one semantic component. The decomposition is based on the cosine similarity between the embedding and the centroid of the cluster, which is the representation of the semantic component. For an embedding $\mathbf{x}$ of a sample $s$, e define the $i$-th residual in an inductive way:
$\mathbf{x}^{\prime}_{0}=\mathbf{x},\\ \mathbf{x}^{\prime}_{i+1}=\mathbf{x}^{\prime}_{i}-\mu\mathbb{1}_{\alpha_{i,i}>\alpha}\langle\mathbf{x}_{i},\mathbf{v}_{i}\rangle\mathbf{v}_{i}$ (3)
where $\mathbf{v}_{i}$ is the $i$-th component and $\alpha_{i,i}=\frac{\langle\mathbf{x}_{i},\mathbf{v}_{i}\rangle}{||\mathbf{x}_{i}||}$. The activation of the sample $\mathbf{x}$ along the $j$-th is thus defined as the dot product of the residual and the component:
$a_{j}=\mu\mathbb{1}_{\alpha_{i,i}>\alpha}\langle\mathbf{x}^{\prime}_{j},\mathbf{v}_{j}\rangle$ (4)
From this, we can assign additional components to the samples by computing the activation along the components and applying a threshold per iteration to choose the relevant ones. This ensures, that the decomposition is able to capture multiple semantic components per sample even beyond the cluster assignments.
8.3 Stopping Criteria
Without additional stopping criteria, the algorithm would spend a large number of iterations decomposing residual embeddings that are already close to zero and lack any meaningful information. To avoid this, we introduce a number of stopping criteria that can we use to terminate the algorithm. These criteria are as follows:
  • F: Terminate after a fixed number of iterations $I$
  • NC-S: Number of new clusters found in the last $S$ iterations is lower than a threshold $T$
  • RN: Matrix 2-norm of residual embeddings lower than a threshold $M$
For our evaluation, we set $I=10$, $S=2$, $T=5$, and $M=0.01$ for all use cases.
9 Extended Results
9.1 Implementation and Preprocessing
The datasets we use for evaluating our method have been obtained from different sources. The Tweets made by Donald Trump are available in the Trump Twitter Archive Brown (2021), from where we download the dataset directly as .csv which resulted in 60,053 samples as of 10th of December 2023. The Chinese News Tweets have been scraped from a list of Chinese news outlets, which we compiled by two criteria: (1) The majority of the Tweets uses simplified or traditional Chinese characters and (2) the news outlet is based outside of the borders of the People’s Republic of China. The list is far from complete, but we stopped adding new accounts when a threshold of 1M Tweets was reached. The accounts included are: “voachinese", “chinesewsj", “abcchinese", “reuterschina", “dw_chinese", “rfi_tradcn", “asiafinance", “initiumnews", “cdtchinese", “kbschinese", “rijingzhongwen", “weiquanwang", “zaobaosg", “rfa_chinese", “rfi_cn", “rcizhongwen", “kbschines", and “chosunchinese".
From this list, we scraped all the Tweets available through the Twitter API which resulted in 1,463,150 samples until the closure of our academic API access in April 2023. The Hausa Tweets have been collected for sentiment classification by Muhammad et al. (2022) and, at the time of acquiring the dataset, amounted to 511,523 samples in Hausa. The current version, of which our sample is a subset of, with slightly more samples (516,523) is available on GitHub. [^3] https://github.com/hausanlp/NaijaSenti/blob/main/sections/unlabeled_twitter_corpus.md. The hyperparameters chosen for each run of the are listed in Figure 4.
For computing the embeddings of the datasets, we used one NVidia A100 GPU, which processed all of the data in less than an hour. The computations for SCA were entirely CPU-based (AMD EPYC 7662) with an allocation of 8 CPU cores and 64 GB of RAM for the Hausa and Chinese News datasets and 16GB for the Trump dataset. On that configuration, the total runtime for SCA (excluding the time to compute the embeddings) on the Trump dataset is 9 minutes, on the Chinese News dataset 345 minutes and on the Hausa dataset 91 minutes. The main bottleneck for the computations is UMAP, of which there exists a GPU implementation through cuML [^4] https://docs.rapids.ai/api/cuml/stable/execution_device_interoperability/. We tested this version and can indeed confirm that the runtimes are much faster using this implementation. However, we found the results to be inconsistent with the CPU version, which is why we decided to stick to the CPU version for the evaluation.
Before embedding the documents, we preprocess the data by removing URLs, mentions, hashtags, emojis, and special characters. The sentence transformer models we use, namely paraphrase-multilingual-mpnet-base-v2 Reimers and Gurevych (2020) for the Trump and Chinese News Tweets and LaBSE Feng et al. (2022), are trained on a large variety of domains and languages. We use the default settings and tokenizers for the models, which are able to handle sequences of up to 512 tokens. We use the umap-learn implementation of UMAP McInnes et al. (2020) with the default settings for dimensionality reduction. For clustering, we use the hdbscan implementation of HDBSCAN Campello et al. (2015) including min_cluster_size and min_samples as hyperparameters of SCA. We use the OCTIS implementation of the coherence measures Terragni et al. (2021).
9.2 Evaluation Metrics
For each run, we calculate a number of statistics to evaluate the performance of the method. We provide the results for the three datasets in the following tables. The statistics are as follows:
  • Noise rate: The fraction of documents that are not assigned to any cluster in the first iteration (1st) and over all iterations.
    $\mathcal{L}_{noise}=\frac{|\{c\in C|c=-1\}|}{|C|}$ (5)
  • Number of clusters: The number of unique clusters in the clustering result.
    $\mathcal{L}_{num}=|C|$ (6)
  • Top 10 topic diversity: As defined in Terragni et al. (2021), it is the number of unique Top-10 representative tokens (as computed by c-TF-IDF weighting) divided by the total number of representative tokens.
Additionaly, we define the sample overlap between sets $C_{1}=\{s_{1},s_{2},...,s_{n}\}$ and $C_{2}=\{s^{\prime}_{1},s^{\prime}_{2},...,s^{\prime}_{m}\}$ as the fraction of elements that are in both clusters.
$o(C_{1},C_{2})=\frac{|C_{1}\cap C_{2}|}{|C_{1}\cup C_{2}|}$ (7)
from this, we define additional metrics to quantify the novelty of the semantic components uncovered by SCA compared to the topics found by BERTopic:
  • the average maximum sample overlap of the clusters found after the first iteration with any of the clusters in the complete HDBSCAN-hierarchy of the first iteration. For $C_{I}=\{C_{1},C_{2},...,C_{M}\}$ the set of possible clusters in the first iteration as defined by the HDBSCAN-hierarchy and $C_{+}=\{C^{+}_{1},C^{+}_{2},...,C^{+}_{N}\}$ the set of clusters found by SCA after the first iteratiom, we define the average maximum sample overlap as:
    $\mathcal{L}_{ASO}=\frac{1}{N}\sum_{i=1}^{N}\max_{j=1,...,M}o(C^{+}_{i},C_{j})$ (8)
    In other words, this metric tries to understand to what extend the additional components found by SCA where present in the topic-hierarchy of BERTopic. High values close to one would indicate, that most compnents found by SCA are sub- or supertopics of the topics found by BERTopic, low values close to zero would indicate that the components found by SCA are not present in the topic-hierarchy of BERTopic and hence provide additional insights.
  • In an analog way, for the component token representations of the first iteratiom $R_{I}=\{R_{1},R_{2},...,R_{M}\}$ and the representations of following iterations $R_{+}=\{R^{+}_{1},R^{+}_{2},...,R^{+}_{N}\}$, we define the average maximum token overlap as:
    $\mathcal{L}_{ATO}=\frac{1}{N}\sum_{i=1}^{N}\max_{j=1,...,M}o(R^{+}_{i},R_{j})$ (9)
The coherence scores are calculated over the token representations of each topic using the respective OCTIS implementations. For some confirmation measure $m:V\times V\rightarrow\mathbb{R}$ that maps two words to a real number, the topic coherence score is defined as
$\mathrm{TC}_{m}=\frac{1}{T}\sum_{i=1}^{T}\frac{1}{\binom{m}{2}}\sum_{r=2}^{K}\sum{s=1}{r-1}m(w^{r}_{i},w^{s}_{i})$ (10)
  • NPMI Coherence Aletras and Stevenson (2013) propose to use the NPMI measure Bouma (2009) due to it being closer to human topic ratings than previous versions.
    $m(w^{r}_{i},w^{s}_{i})=NPMI(w^{r}_{i},w^{s}_{i})=\frac{\log_{2}\frac{p(w^{r}_{i},w^{s}_{i})+\epsilon}{p(w^{r}_{i})p(w^{s}_{i})}}{-\log_{2}(p(w^{r}_{i},w^{s}_{i})+\epsilon)}$ (11)
  • CV Coherence Röder et al. (2015) propose to use the CV measure, which scales the coherence terms of NPMI by a power of a parameter $\gamma$.
    $m(w^{r}_{i},w^{s}_{i})=\mathrm{CV}(w^{r}_{i},w^{s}_{i})=\mathrm{NPMI}^{\gamma}(w^{r}_{i},w^{s}_{i})$ (12)
Statistic D.T. Hau. News
$\alpha$ 0.20 0.10 0.10
$\mu$ 0.95 1.00 1.00
min_cluster_size 100 300 300
min_samples 50 300 300
Overlap Threshold $\theta$ 0.5 0.5 0.5
No. of Components (1st) 55 102 212
No. of Components 182 331 594
No. of Clusters 190 340 613
Noise Rate (1st) 0.275 0.574 0.429
Noise Rate 0.000 0.001 0.000
Avg. Maximum Sample Overlap 0.079 0.093 0.077
Avg. Maximum Token Overlap 1.387 1.852 1.318
NPMI Coherence -0.168 -0.061 0.120
CV Coherence 0.360 0.400 0.595
Topic Diversity 0.703 0.804 0.803
NPMI Coherence (1st) -0.135 -0.028 0.186
CV Coherence (1st) 0.370 0.431 0.683
Topic Div. (1st) 0.947 0.909 0.929
Figure 4: Run statistics for the three datasets. (1st) denotes the statistic calculated only over the clusters/components of the first iteration of SCA, which is equivalent to a BERTopic run with the same hyperparameters.
No. of components
$\mu$/$\alpha$ 0.0 0.05 0.1 0.2 0.3 0.4 0.5
0.5 558 416 384 218 195 577 945
0.6 582 421 410 285 208 303 680
0.7 491 444 364 305 107 107 177
0.8 363 341 358 317 281 252 108
0.9 466 365 264 319 292 246 371
0.95 336 353 347 336 231 250 302
1.0 333 358 353 336 287 289 261
Topic diversity
$\mu$/$\alpha$ 0.0 0.05 0.1 0.2 0.3 0.4 0.5
0.5 0.539 0.631 0.635 0.728 0.710 0.259 0.096
0.6 0.530 0.651 0.663 0.690 0.726 0.553 0.213
0.7 0.611 0.643 0.697 0.699 0.859 0.857 0.709
0.8 0.702 0.702 0.706 0.725 0.706 0.689 0.853
0.9 0.655 0.713 0.749 0.714 0.713 0.710 0.589
0.95 0.728 0.710 0.724 0.724 0.746 0.705 0.638
1.0 0.725 0.719 0.716 0.717 0.720 0.682 0.658
Noise rate
$\mu$/$\alpha$ 0.0 0.05 0.1 0.2 0.3 0.4 0.5
0.5 0.003 0.003 0.004 0.000 0.000 0.221 0.398
0.6 0.001 0.003 0.010 0.012 0.003 0.001 0.240
0.7 0.002 0.012 0.001 0.004 0.000 0.000 0.000
0.8 0.000 0.002 0.001 0.001 0.000 0.016 0.001
0.9 0.000 0.000 0.005 0.004 0.000 0.000 0.000
0.95 0.000 0.006 0.001 0.000 0.000 0.000 0.001
1.0 0.002 0.003 0.003 0.001 0.000 0.000 0.003
NPMI Coherence
$\mu$/$\alpha$ 0.0 0.05 0.1 0.2 0.3 0.4 0.5
0.5 -0.192 -0.191 -0.187 -0.162 -0.160 -0.169 -0.196
0.6 -0.202 -0.198 -0.198 -0.178 -0.161 -0.168 -0.173
0.7 -0.202 -0.203 -0.195 -0.186 -0.153 -0.153 -0.150
0.8 -0.191 -0.197 -0.197 -0.190 -0.182 -0.177 -0.152
0.9 -0.206 -0.198 -0.194 -0.194 -0.187 -0.180 -0.197
0.95 -0.199 -0.199 -0.198 -0.197 -0.180 -0.179 -0.188
1.0 -0.200 -0.201 -0.200 -0.201 -0.191 -0.193 -0.181
CV Coherence
$\mu$/$\alpha$ 0.0 0.05 0.1 0.2 0.3 0.4 0.5
0.5 0.356 0.362 0.362 0.368 0.368 0.364 0.353
0.6 0.355 0.361 0.357 0.363 0.366 0.372 0.359
0.7 0.361 0.363 0.358 0.365 0.374 0.374 0.374
0.8 0.359 0.365 0.362 0.365 0.361 0.366 0.375
0.9 0.359 0.358 0.367 0.363 0.361 0.363 0.359
0.95 0.365 0.364 0.362 0.359 0.362 0.367 0.360
1.0 0.369 0.361 0.362 0.365 0.360 0.365 0.365
Figure 5: Grid run results for different hyperparameters. All configurations were run with $\theta=1.0$ to avoid occluding “bad" results by merging many components.
9.3 Trump Dataset: Components
ID N c-TF-IDF representation and medoid
0 14257 apprentice, celebapprentice, celebrity, interviewed, mr, run, season, ratings, trump2016, please
"""@djf123: @realDonaldTrump @PJTV #DonaldWillBeTheGOPNominee!"
1 1634 entrepreneurs, yourself, success, focus, champion, passion, negotiation, learn, art, momentum
“A very good way to pave your own way to success is simply to work hard and to be diligent” – Think Like a Champion
2 1217 rally, crowd, join, tickets, crowds, maga, south, evening, landed, soon
Leaving Michigan now, great visit, heading for Iowa. Big Rally!
3 1215 border, wall, immigration, southern, illegal, laws, immigrants, patrol, drugs, borders
Our Southern Border is under siege. Congress must act now to change our weak and ineffective immigration laws. Must build a Wall. Mexico, which has a massive crime problem, is doing little to help!
4 936 sleepy, fracking, 47, suburbs, firefighter, pack, basement, radical, abolish, wins
A vote for Joe Biden is a vote to extinguish and eradicate your state’s auto industry. Biden is a corrupt politician who SOLD OUT Michigan to CHINA. Biden is the living embodiment of the decrepit and depraved political class that got rich bleeding America Dry!
5 893 hotel, chicago, tower, luxury, building, rooms, restaurant, hotels, sign, spa
.@TrumpSoHo features a striking glass walled building w/ loft inspired interiors http://t.co/yUn5d14t0Q NYC’s trendiest luxury hotel
6 891 china, tariffs, trade, currency, products, tariff, countries, barriers, goods, billions
….with the U.S., and wishes it had not broken the original deal in the first place. In the meantime, we are receiving Billions of Dollars in Tariffs from China, with possibly much more to come. These Tariffs are paid for by China devaluing, pumping, not by the U.S. taxpayer!
7 852 investigation, report, dossier, director, spy, emails, documents, collusion, page, obstruction
The Mueller probe should never have been started in that there was no collusion and there was no crime. It was based on fraudulent activities and a Fake Dossier paid for by Crooked Hillary and the DNC, and improperly used in FISA COURT for surveillance of my campaign. WITCH HUNT!
8 838 discussing, obama, records, college, birth, charity, interview, certificate, applications, offer
What is your thought as to why Obama refused millions for charity and did not show his records and applications?
9 702 golf, course, links, courses, club, monster, blue, championship, international, aberdeen
Named best golf course in the world by @RobbReport, Trump Int’l Golf Links Scotland is a 7,400 yd par 72 http://t.co/7FYb3dEQcD
Figure 6: Iteration 1: Top-10 components by number of samples N together with their c-TF-IDF representations and medoids.
ID N c-TF-IDF representation and medoid Overl.
55 2468 run, please, needs, 2016, need, runs, trump2016, trumpforpresident, running, mr 0.4
"""@tlrchrstphrbrsn: @realDonaldTrump please run for president""" 0.75
56 2099 fake, media, nytimes, news, failing, sources, story, dishonest, reporting, stories 0.0
So much Fake News being put in dying magazines and newspapers. Only place worse may be @NBCNews, @CBSNews, @ABC and @CNN. Fiction writers! 0.54
57 984 apprentice, celebrity, season, celebrityapprentice, celebapprentice, episode, nbc, cast, wait, 13th 0.4
Celebrity Apprentice is rebroadcasting last weeks episode at 9 P.M. WITH A GREAT NEW EPISODE FEATURING @MELANIA TRUMP AT 10 P.M. - AMAZING! 0.70
58 767 inspiration, model, tweets, role, thanks, entrepreneur, genius, inspiring, tweet, birthday 0.0
"""@balango212 Donald J. Trump, my biggest inspirational in life."" Thank you." 0.49
60 625 condolences, prayers, deepest, heroes, welcome, victims, families, honor, thoughts, shooting 0.3
Saddened to hear the news of civil rights hero John Lewis passing. Melania and I send our prayers to he and his family. 0.22
61 604 truth, hate, honesty, haters, speaks, afraid, speak, honest, losers, listen 0.0
"""@HeyMACCC: I feel like the only people who hate @realDonaldTrump are people who can’t handle the truth…he tells it how it is.""" 0.27
62 598 oscars, ratings, boring, dumbest, show, dummy, overrated, tv, television, comedian 0.1
@DanielRobinsMD. @Lawrence is a very dumb guy who just doesn’t know it-his show is a critical and ratings disaster! 0.40
63 539 hunt, witch, collusion, obstruction, conflicted, investigation, evidence, report, angry, witnesses 0.4
The Greatest Witch Hunt In American History! 0.27
64 534 approval, rating, poll, leads, overall, surges, party, republican, 52, polls 0.1
96% Approval Rating in the Republican Party. 51% Approval Rating overall in the Rasmussen Poll. Thank you! 0.20
65 501 entrepreneurs, touch, yourself, focus, goals, opportunities, momentum, passion, vision, requires 0.5
Entrepreneurs: Stay focused and be tenacious. Pay attention to people who know what they’re talking about. Stay fixed on your goals! 0.77
Figure 7: Iteration 2: Top-10 components by number of samples N together with their c-TF-IDF representations and medoids. The overlap scores refer to the maximum token overlap (top) with any component from the first iteration, i.e. a BERTopic run, and the maximum sample overlap (bottom) with any cluster in the cluster hierarchy of HDBSCAN from the first iteration (i.e. all possible topics that could be found through BERTopic).
ID N c-TF-IDF representation and medoid Overl.
132 1502 hotel, golf, course, tower, luxury, links, club, resort, hotels, blue 0.5
We do have a Trade Deficit with Canada, as we do with almost all countries (some of them massive). P.M. Justin Trudeau of Canada, a very good guy, doesn’t like saying that Canada has a Surplus vs. the U.S.(negotiating), but they do…they almost all do…and that’s how I know! 0.52
135 554 hurricane, prayers, responders, condolences, storm, victims, coast, local, water, rescue 0.4
I have instructed the United States Navy to shoot down and destroy any and all Iranian gunboats if they harass our ships at sea. 0.27
137 470 ratings, fox, biased, credibility, daytime, anchors, network, debates, networks, loser 0.1
RT @DailyCaller: CNN Receives Bad TV Rating For Speaker Pelosi Town Hall https://t.co/YxzzHljJAK 0.19
138 440 failing, apologize, nytimes, magazine, apology, apologized, times, paper, circulation, newspaper 0.0
The New York Times has apologized for the terrible Anti-Semitic Cartoon, but they haven’t apologized to me for this or all of the Fake and Corrupt news they print on a daily basis. They have reached the lowest level of “journalism,” and certainly a low point in @nytimes history! 0.22
140 424 sources, enemy, fake, lie, hoax, lied, lies, media, stories, false 0.1
"Anytime you see a story about me or my campaign saying ""sources said,"" DO NOT believe it. There are no sources, they are just made up lies!" 0.17
141 351 economy, change, changing, stronger, quarter, market, stock, growth, booming, wand 0.2
Our economy is perhaps BETTER than it has ever been. Companies doing really well, and moving back to America, and jobs numbers are the best in 44 years. 0.13
142 340 leaving, crowd, landed, heading, evening, landing, soon, south, ready, departing 0.5
I’m leaving now for Ireland, Spain, Scotland and elsewhere–crazy life! 0.17
143 338 billionaire, businessman, smartest, business, rich, genius, smart, chess, investor, businessmen 0.0
"""It’s more important to be smart than tough. I know businessmen who are brutally tough, but they’re not smart.” – Think Like A Billionaire" 0.05
145 300 rather, wish, prefer, would, better, awesome, actor, achieve, greatest, best 0.1
"""@iamapatsfan: I would rather Donald Trump be the president than Jeb Bush."" Would do a much better job!" 0.05
146 280 penalty, killer, prison, death, prisoners, hostage, hostages, jail, sentenced, thug 0.0
caught, he cried like a baby and begged for forgiveness…and now he is judge, jury. He should be the one who is investigated for his acts. 0.04
Figure 8: Iteration 4: Top-10 components by number of samples N together with their c-TF-IDF representations and medoids. The overlap scores refer to the maximum token overlap (top) with any component from the first iteration, i.e. a BERTopic run, and the maximum sample overlap (bottom) with any cluster in the cluster hierarchy of HDBSCAN from the first iteration (i.e. all possible topics that could be found through BERTopic).
9.4 Chinese News Outlet Tweets: Components
Figure 9: Iteration 1: Top-10 components by number of samples N together with their c-TF-IDF representations and medoids.
Figure 10: Iteration 2: Top-10 components by number of samples N together with their c-TF-IDF representations and medoids. Overlap describes the maximum number of tokens a component’s representation has in common with any component from the first iteration.
Figure 11: Iteration 3: Top-10 components by number of samples N together with their c-TF-IDF representations and medoids. Overlap describes the maximum number of tokens a component’s representation has in common with any component from the first iteration.
9.5 Hausa Dataset: Components
ID N c-TF-IDF representation and medoid with (GPT-4o translations)
0 10509 masha, insha, yasaka, allahu, jikansa, akbar, yajikansa, jikanshi, rahamarsa, jikan
blessing, will, you, God, your child, great, your child, your child, our mercy, child
afwan da allah wayece rahama sadau
May Allah forgive and have mercy on us.
1 10185 musulunci, musulmai, musulmi, addinin, musulma, musulinci, islam, addini, muslim, islama
Islam, Muslims, Muslims, religion, Muslim, Islam, Islam, religion, Muslim, Islam
karya kike dama tun fari kintaba ridda wanna tufafi da kikasa sukuma namusulmaine musulumci yayadda da tabarrunne keda kanki kyasan bakida alaka da musulumci
You are lying, you have worn this outfit before and it is not Islamic. Islam does not agree with showing off, and you know you have no connection with Islam.
2 7341 madrid, barcelona, liverpool, chelsea, arsenal, manchester, real, gasar, league, united
Madrid, Barcelona, Liverpool, Chelsea, Arsenal, Manchester, Real, Gasar, League, United
premier league an kama mai kalaman wariya a wasan manchester
Premier League arrests a player for racist remarks in the Manchester match.
3 7161 coronavirus, covid, corona, virus, rigakafin, korona, kamu, cutar, allurar, kamuwa
coronavirus, covid, corona, virus, vaccine, corona, infection, disease, vaccine, infection
jumillar maaikatan lafiya sun kamu da cutar coronavirus a kano
Many people in Kano have contracted the coronavirus.
4 6244 bindiga, garkuwa, hallaka, rundunar, sanda, cafke, gobara, sandan, harin, arin
gun, shield, destroy, army, police, arrest, fire, policeman, attack, attack
yan bindiga sun yi garkuwa da mutum a shirori jihar neja
Gunmen have kidnapped a person in Shiroro, Niger State.
5 5400 bbc, bbchausa, hausa, news, tarihin, kubani, sashen, labarai, bakuda, haussa
bbc, bbchausa, hausa, news, history, give me, section, news, you don’t have, hausa
kai bbc makaryatanebasu gayan aibin kasarsu senawasu
BBC is not giving them the opportunity to speak about their
6 4769 gaskiyane, gaskiyar, gaske, gaskiya, true, gaskiyan, maganarka, adalci, adalcin, halinta
truth, truth, true, truth, true, truths, your statement, justice, justice, your character
gaskiya dai abun yfara katsanta
Truly, the situation is becoming
7 4672 nigerian, niger, nigeria, nigeriya, nigeriar, talakan, kasata, nigerians, talakawan, republic
Nigerian, Niger, Nigeria, Nigeria, Nigeria, poverty, wealth, Nigerians, poor people, republic
wae ahakan ake so acanza nigeria ana irin wanan zalincin karara
Why do they want to change Nigeria with this kind of blatant oppression?
8 4167 hmmm, hmm, hmmmm, hmmmmm, hm, uhmm, uhm, uhmmm, hmmmmmm, hummm
hmmm, hmm, hmmmm, hmmmmm, hm, uhmm, uhm, uhmmm, hmmmmmm, hummm
hmmmm tohhh dan masani jikan masanawa ae sae kaje ka sakosu kawae
Hmmm, well, an expert knows the offspring of experts, so just go and ask them.
9 3953 yarjejeniyar, israila, bukaci, syria, iran, turkiya, majalisar, dokar, sudan, fadarsa
diplomacy, Israel, request, Syria, Iran, Turkey, council, law, Sudan, his/her place
majalisar najeriya ta bukaci buhari ya kafa dokartabaci
The Nigerian parliament has asked Buhari to declare a state of emergency.
Figure 12: Iteration 1: Top-10 components by number of samples N together with their c-TF-IDF representations and medoids.
ID N c-TF-IDF representation and medoid with (GPT-4o translations) Overl.
102 19015 shiru, magana, gaya, kanku, maganar, fada, said, naka, miki, kike 0.0
silent, talk, tell, yourselves, the talk, say, said, yours, to you, you are
kaji wata magana ta banxa da kake fada wadane manya kasa ku kenan kune manya kasa dole kace hakan tunda abin bashi kai kanka ba
Understand a useless talk that you are saying, you elders of the land, you are the elders of the land, you must say that since the matter does not concern you.
103 6995 allahumma, hai, summa, makomarsa, jikanshi, jikansa, subhanallah, makoma, innaka, yajikansa 0.3
Oh Allah, yes, then, his place, his body, his body, glory be to Allah, place, indeed you, his body
assalamu alekum barkammudawarhaka
Peace be upon you
104 6212 yasan, san, gane, sani, fahimci, sanin, fahimta, fahimtar, yasani, sane 0.0
know, know, understand, know, understand, knowing, understanding, understanding, knows, aware
wanda yasan mene kano shine xai fahimci abinda jagoran arewa yake nupi sannan ya gane nupinsa
Whoever understands what the northern leader is saying will also understand his own language.
105 6077 gani, kallo, see, kallon, ganin, tunanin, ganina, duba, samodara, kalli 0.1
look, see, see, seeing, seeing, thinking, my sight, check, samodara, look
rahma sadau kinyi gaba mahasada sai kallo magulmata sai gani
Rahma Sadau, keep moving forward, let the envious watch, and let
106 4846 matsala, talauci, wahala, matsalar, matsalan, matsaloli, matsalolin, kalubale, problem, wuya 0.1
problem, poverty, trouble, problem, problem, problems, problems, challenge, problem, difficulty
ya zamto cikin manyan matsalolin kasar da yana da wahalar samuwaidan kuma ya samu ga yawan alala da karancin nagarta matukar basu nemo bakinzaren ba to dole su kasance cikin wahala da rudani yau an wayi gari matsalar ta shiga gidan kowa
We are facing major problems in the country that are difficult to address, and with the abundance of challenges and lack of quality, if they do not find solutions, they will inevitably remain in hardship and confusion; today, the problem has entered everyone’s home.
107 4809 kotu, hukunci, doka, hukuncin, kotun, hukunta, yanke, sharia, koli, dokar 0.1
court, judgment, law, judgment, the court, to judge, to decide, Islamic law, large, the law
wannan hukunci ne bisa doran doka wannda ya ga bai yi daidai ba to se ya daukaka kara
This is a judgment based on the law, and whoever feels it is not right should appeal.
108 4594 dawo, koma, dawowa, maimaita, mayar, komawa, back, sake, maida, mashekiya 0.1
return, go back, return, repeat, give back, return, back, redo, return, return
tabbijanledoci sai kasata nageriya
The government will not allow Nigeria
109 4141 jikansu, tona, kubutar, garesu, yabasu, asirinsu, maganinsu, taimakesu, gafarta, sharrinsu 0.2
sacrifice, light, to save, to help, to speak, their secrets, their words, to assist, to forgive, their evil
qoqari su kayi da zaa jinjina musu
Efforts that deserve commendation.
110 3879 aikin, aiki, aikinka, aikinku, aikinsa, office, business, job, aikinsu, work 0.0
work, work, your work, your work, his work, ofis, kasuwanci, aiki, their work, work
aiki aiki aiki wani aikin sai kano
Work, work, work, what work except Kano.
111 3779 musulmai, musulunci, musulmi, islam, musulinci, addinin, musulman, musulmin, khadimul, musulunchi 0.6
muslims, islam, muslim, islam, islam, religion, muslim, muslims, servant, muslim
Muslim, Islam, Muslim, Islam, Islam, religion, Muslim, Muslim, servant, Muslim
komi kafiri xaiyi makiyin musulunci ne da musulmai
The infidels are the enemies of Muslims.
Figure 13: Iteration 2: Top-10 components by number of samples N together with their c-TF-IDF representations and medoids. Overlap describes the maximum number of tokens a component’s representation has in common with any component from the first iteration.
ID N c-TF-IDF representation and medoid with (GPT-4o translations) Overl.
253 13223 talauci, kidnappers, talaka, yunwa, talakawa, bandits, sannan, rashin, kidnapping, tausayin 0.2
suffering, kidnappers, poor person, hunger, poor people, bandits, then, lack, kidnapping, compassion
daman munsani kuma mundaina binsa
We have seen the light and we have
254 7468 tuba, manzon, kika, miki, kiyi, kiji, saw, yafe, w, kikayi 0.2
repent, messenger, you did, to you, do, you know, see, forgive, and, you did
kaga maiyi don allah kayi naka kayina rogo malam munan daram muna tayaka daaddua allah kareka daga sharrin makiaya malam inda tun abaya gwamanati natafiya da ireirenku da sai allah yasan matakin damuka taka
Please, for the sake of Allah, do your part and help us, we are in need of your assistance. May Allah protect you from the evil of the oppressors. Since the time of the government, you have been like this, and only Allah knows the extent of your actions.
255 7068 babu, zamuce, bace, bazan, no, baiyi, taba, dace, daceba, bazai 0.0
father, we will, lost, I will not, no, he/she/it did not, touch, good, it is good, he/she/it will not
babu wanda zai yayemuna wannan sai allah addua kawai zamuyi janaa
There is no one who can do this for us except God, we will only pray.
256 6758 kina, kike, kanku, ruwanku, kunyi, kika, zaki, uwarku, kukeyi, baku 0.0
your, you, your, your water, you, you, lion, your mother, you are doing, not you
kana shugaba bakasan me zakai
You don’t know what you will do.
257 5190 zaiyi, zasuyi, zatayi, zamuyi, will, insha, bige, sabuwa, chacha, zaiga 0.1
we will, they will do, they will sell, we will do, will, if God wills, beat, new, new, he will go
najeriya zata lallasa afirka takudu
Nigeria will dominate Africa.
258 4818 koshin, acikin, lauje, nadi, yantattun, shiga, mati, muje, fckng, keep 0.1
outside, inside, leaf, pound, free, enter, mat, let’s go, fckng, keep
to kai meye ruwanka a ciki munafuki
Why are you involving yourself in this, hypocrite?
259 4622 ronaldo, league, tottenham, united, psg, juventus, mourinho, messi, bayern, champions 0.2
Ronaldo, liga, Tottenham, United, PSG, Juventus, Mourinho, Messi, Bayern, şampiyonlar
kasuwar cinikin yan kwallon kafa makomar ronaldo sterling jesus nunez romero giroud
football transfer market future of Ronaldo Sterling Jesus Nunez Romero Giroud
260 3477 bindiga, garkuwa, cafke, neja, sanda, karamar, rundunar, sandan, hallaka, hatsarin 0.7
gun, shield, capture, snake, fence, small, battalion, soldier, ambush, accident
wani dan bindiga ya kai hari wata makaranta a birnin kazan da ke kasar rasha da safiyar ranar talata ya kashe mutum takwas bakwai daliban da ke aji takwas da kuma malami guda ya kuma ji wa wasu raunukan da suka kai su ga kwanciya a asibiti a cewar jamian rasha ap
A gunman attacked a school in the city of Kazan, Russia, on Tuesday morning, killing eight people, seven of whom were eighth-grade students and one teacher, and injuring others who were hospitalized, according to Russian officials.
261 2797 congratulations, hai, allahumma, main, mein, congrats, innaka, pyaar, tum, hoon 0.2
congratulations, hi, O Allah, I, in, congratulations, indeed, love, you, am
fakuluma kaddararrahmanu mafulunwamaiyatawakkalu alallahu fahuwa hasbuk
Speak about the destiny of the Most Merciful
262 2651 said, gwara, sunce, gaya, kauye, fada, bahaushe, nace, makaho, aniya 0.0
said, told, they said, tell, village, say, Hausa person, I said, blind person, intention
kauye garin su kare yace nima kaini akai kaidai ai shaani danbirni yacuci na kauye
The village said take me there too, but the city deceived the villager.
Figure 14: Iteration 3: Top-10 components by number of samples N together with their c-TF-IDF representations and medoids. Overlap describes the maximum number of tokens a component’s representation has in common with any component from the first iteration.
9.6 Performance of Transformed Embeddings on MTEB
We benchmark the performance of the embeddings generated by a pipeline of SBERT and the SCA transformation against the original SBERT embeddings and an embedding consisting of the first 190 principal components found through PCA, which is equal to the number of components found by SCA trained on the full Trump dataset. Furthermore, we include the same benchmark on components from just the first iteration and a 55-dimensional PCA.
As a linear decomposition, SCA is a transformation of the embeddings into an explainable space, where each dimension corresponds to the score along one semantic component as defined by formula (2) (see Appendix 8.2 for details). We use this interpretation of the component scores to evaluate the topic distribution in terms of its performance as a representation of samples in downstream tasks defined in the MTEB benchmark (Muennighoff et al., 2023) and compare against the vanilla embeddings as well as a PCA transformation of the same dimension.
The main results of all chosen benchmarks are shown in table 15. SCA as well as PCA perform uniformly well over all benchmarks and match the performance of the base model with minor variations.
task name type metric SBERT SCA PCA SCA’ PCA’
AmazonCounterfactualClassification Class. accuracy 0.758 0.752 0.757 0.703 0.737
AmazonPolarityClassification Class. accuracy 0.764 0.768 0.768 0.778 0.781
AmazonReviewsClassification Class. accuracy 0.385 0.385 0.385 0.370 0.390
ArguAna Retr. ndcg_at_10 0.489 0.486 0.479 0.456 0.457
ArxivClusteringP2P Clust. v_measure 0.379 0.371 0.374 0.352 0.350
ArxivClusteringS2S Clust. v_measure 0.317 0.314 0.315 0.299 0.302
AskUbuntuDupQuestions Rerank. map 0.602 0.609 0.606 0.595 0.599
BIOSSES STS spearman 0.763 0.754 0.765 0.711 0.750
Banking77Classification Class. accuracy 0.811 0.802 0.808 0.768 0.789
BiorxivClusteringP2P Clust. v_measure 0.330 0.326 0.318 0.291 0.287
BiorxivClusteringS2S Clust. v_measure 0.294 0.292 0.292 0.270 0.269
CQADupstackRetrieval Retr. ndcg_at_10 0.313 0.309 0.314 0.268 0.278
ClimateFEVER Retr. ndcg_at_10 0.153 0.151 0.156 0.132 0.126
DBPedia Retr. ndcg_at_10 0.262 0.259 0.262 0.217 0.226
EmotionClassification Class. accuracy 0.459 0.454 0.457 0.438 0.440
FEVER Retr. ndcg_at_10 0.568 0.553 0.553 0.469 0.458
FiQA2018 Retr. ndcg_at_10 0.230 0.224 0.231 0.183 0.194
HotpotQA Retr. ndcg_at_10 0.370 0.350 0.372 0.248 0.263
ImdbClassification Class. accuracy 0.646 0.647 0.646 0.647 0.644
MSMARCO Retr. ndcg_at_10 0.571 0.555 0.568 0.537 0.541
MTOPDomainClassification Class. accuracy 0.892 0.880 0.887 0.839 0.863
MTOPIntentClassification Class. accuracy 0.687 0.658 0.675 0.595 0.632
MassiveIntentClassification Class. accuracy 0.693 0.684 0.688 0.652 0.667
MassiveScenarioClassification Class. accuracy 0.762 0.749 0.761 0.715 0.741
MedrxivClusteringP2P Clust. v_measure 0.319 0.317 0.320 0.305 0.307
MedrxivClusteringS2S Clust. v_measure 0.315 0.311 0.317 0.305 0.306
MindSmallReranking Rerank. map 0.302 0.300 0.292 0.301 0.289
NFCorpus Retr. ndcg_at_10 0.255 0.257 0.259 0.230 0.234
NQ Retr. ndcg_at_10 0.336 0.329 0.340 0.277 0.306
QuoraRetrieval Retr. ndcg_at_10 0.864 0.862 0.863 0.850 0.853
RedditClustering Clust. v_measure 0.457 0.455 0.457 0.414 0.416
RedditClusteringP2P Clust. v_measure 0.521 0.513 0.516 0.481 0.484
SCIDOCS Retr. ndcg_at_10 0.140 0.140 0.141 0.122 0.125
SICK-R STS spearman 0.796 0.793 0.798 0.783 0.787
STS12 STS spearman 0.779 0.784 0.759 0.785 0.755
STS13 STS spearman 0.851 0.849 0.852 0.829 0.842
STS14 STS spearman 0.808 0.808 0.797 0.793 0.783
STS15 STS spearman 0.875 0.872 0.868 0.856 0.859
STS16 STS spearman 0.832 0.833 0.820 0.826 0.812
STS17 STS spearman 0.870 0.866 0.858 0.840 0.843
STS22 STS spearman 0.635 0.637 0.608 0.632 0.599
STSBenchmark STS spearman 0.868 0.866 0.858 0.853 0.847
SciDocsRR Rerank. map 0.781 0.778 0.774 0.761 0.754
SciFact Retr. ndcg_at_10 0.503 0.500 0.502 0.429 0.442
SprintDuplicateQuestions PClass. similarity_ap 0.906 0.912 0.911 0.892 0.897
StackExchangeClustering Clust. v_measure 0.530 0.531 0.535 0.493 0.501
StackExchangeClusteringP2P Clust. v_measure 0.331 0.333 0.332 0.328 0.325
StackOverflowDupQuestions Rerank. map 0.468 0.462 0.466 0.449 0.452
SummEval Summ. spearman 0.316 0.321 0.318 0.312 0.314
TRECCOVID Retr. ndcg_at_10 0.379 0.363 0.383 0.345 0.370
Touche2020 Retr. ndcg_at_10 0.174 0.176 0.177 0.168 0.173
ToxicConversationsClassification Class. accuracy 0.656 0.655 0.656 0.656 0.657
TweetSentimentExtractionClassification Class. accuracy 0.592 0.604 0.604 0.595 0.604
TwentyNewsgroupsClustering Clust. v_measure 0.444 0.449 0.450 0.434 0.419
TwitterSemEval2015 PClass. similarity_ap 0.667 0.663 0.667 0.647 0.664
TwitterURLCorpus PClass. similarity_ap 0.851 0.847 0.851 0.831 0.844
Average 0.552 0.549 0.550 0.524 0.529
Figure 15: MTEB evaluation results